Whats the cosine rule.
 
 

Grade:12

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
\\c^2=a^2+b^2-2abcosC \\b^2=c^2+a^2-2accosB \\a^2=b^2+c^2-2cbcosA
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty